The price of Dwayne Corporation stock is expected to be RM68 in 5 years. Dividends
are anticipated to increase at an annual rate of 10 percent from the most recent
dividend of RM1.00. If your required rate of return is 15 percent, how much are you
willing to pay for Dwayne stock now?
1
Expert's answer
2015-03-20T10:55:09-0400
The price of Dwayne Corporation stock is expected to be RM68 in 5 years. g = 10%, div0 = RM1.00, r = 15%. The formula for the present value of a stock with constant growth is the estimated dividends to be paid divided by the difference between the required rate of return and the growth rate.
P = div0/(r - g) = 1/(0.15 - 0.1) = RM20.